# DMAligner: Enhancing Image Alignment via Diffusion Model Based View Synthesis

### v1

@@ 1 Introduction
The alignment of consecutive frames is a critical task in computer vision, supporting a wide range of applications such as High Dynamic Range (HDR) imaging [53, 6, 8], video stabilization [26], image deblurring [58], and super-resolution [5]. Precise alignment ensures the accurate merging of information from consecutive frames, leading to enhanced image quality and better performance in subsequent tasks [37, 51, 52]. Despite its significance, the task of image alignment remains challenging due to various factors, including scene dynamics, large camera motion, and varying lighting conditions.
@@
Figure 1: (a) Conventional image alignment based on optical flow and image warping, resulting in ghosting artifacts and occlusion. (b) Our DMAligner directly generate the complete alignment image via diffusion-based view synthesis.
@@
Recent advancements in deep learning have led to substantial progress in neural network-based alignment methods. Generally, image alignment can be divided into three levels: global-level, mid-level, and pixel-level alignment. Global-level alignment typically involves estimating a homography matrix, which can be accomplished using unsupervised deep learning frameworks [23, 25]. This technique effectively aligns images under conditions of planar scenes or when the camera undergoes basic rotational or translational motion. Moreover, through global-to-local homography flow refinement, it can be naturally extended to local mesh-grid homography estimation, allowing it to overcome the limitations of a single homography [23]. However, global- and mid-level methods are restricted in their ability to address more complex scenarios, like parallax, non-rigid deformations, or significant foreground motion, since they assume single or multiple global transformations across the entire image.
@@
Another prominent technique involves the pixel-level approaches with optical flow and image warping [11, 2, 49, 63]. Despite progress in image alignment, current optical flow-based methods face challenges that limit their effectiveness in practical applications. A major issue is the sensitivity of optical flow methods to brightness changes. Variations in lighting conditions, such as shadows, reflections, or changes in ambient light, can significantly distort the brightness constancy assumption, leading to incorrect motion estimation. Besides, the warping process inherent in optical flow-based alignment can introduce various artifacts. Inaccuracies in the motion field can lead to distortions such as stretching, compressing, or tearing of image regions. Additionally, the many-to-one mapping inherent in image warping on occlusion regions inevitably results in ghosting artifacts, as depicted in Fig. 1 (a), even with precise optical flow estimation. Such artifacts not only degrade the visual quality of the aligned frames but also affect the accuracy of subsequent processing tasks.
@@
In this paper, we introduce DMAligner, a diffusion-based framework for image alignment through alignment-oriented view synthesis. As depicted in Fig. 2, the desired image alignment result represents a scenario where the camera pose stays fixed at $P_{1}$ while time advances from $t_{1}$ to $t_{2}$. DMAligner is designed to handle the challenges in image alignment from a new perspective with a generation-based solution (see Fig. 1 (b)), exhibiting strong capabilities and avoiding issues associated with homography and flow-based image warping. Specifically, we first create Dynamic Scene Image Alignment (DSIA) dataset, designed to simulate typical image alignment challenges including camera shifts, moving objects, and fluctuating illumination. Moreover, we propose the Dynamics-aware Diffusion Training approach on the Latent Diffusion Model [38] for learning conditional image generation. The key idea is to enhance the network’s capacity to recognize dynamic information during diffusion model training. This is realized through the Dynamics-aware Mask Producing (DMP) module, which effectively discriminates dynamic foreground regions from static backgrounds within the hidden features across frames. Additionally, we validate the effectiveness of DMAligner through extensive testing on synthetic datasets, including the proposed DSIA and Sintel [3], as well as on the widely-used real-world dataset DAVIS [4]. Both the quantitative and qualitative results demonstrate significant enhancements over contemporary methods. The contributions are summarized as follows:
@@ 2 Related Work
Traditional approaches primarily rely on feature-based methods, where keypoints are detected and matched across images, followed by geometric transformation estimation to align the images [22, 62]. In the deep learning era, DCNNs [68] have revolutionized image alignment in feature space, which is widely applied in video super-resolution [46], video denoising [61], and burst image restoration [10]. In image space, Global- and middel-level alignments typically involve estimating a homography matrix [25] and local mesh-grid homography [23], respectively. But they cannot handle complex scenarios, involving parallax, non-rigid deformations, or significant foreground motion. A notable advancement in image alignment is the adoption of optical flow-based image warping techniques. Optical flow calculates pixel-wise motion between consecutive images, enabling the warping of one image to align with another. With the advances of recent optical flow approaches [36, 41, 29, 7, 33], this method has found widespread use across a variety of video-based applications [27, 1, 2, 63, 11, 55]. Despite their success, warping in occlusion regions inevitably leads to ghosting artifacts, even with accurate flow, as shown in Fig. 1. In contrast, our DMAligner addresses the challenges in image alignment from a new perspective, performing alignment-oriented view synthesis using a diffusion-based framework.
@@
Diffusion models own prominent generative performance for synthesis [56, 14], and various approaches have been developed to improve the performance or efficiency, such as DDPM [14], DDIM [42], LDM [38] and DDNM [50] etc. Different applications have shown good performances when empowered with diffusion models, including but not limited to, optical flow estimation [39, 28], shadow removal [30], image rectangling [67], homography estimation [20], rolling shutter removal [57], low-light image enhancement [15, 16] and camouflaged object detection [65]. In this work, we propose an end-to-end diffusion-based framework that specifically handles cross-frame dynamic information for image alignment.
@@ 3.1 DSIA Dataset Generation
Specifically, inspired by PointOdyssey [66] and Kubric [13], we assemble 25 characters and 100 objects with diverse shapes and materials from several open-source sources, including BlenderKit, Mixamo, GSO [9], and PartNet [35]. These characters are animated using real-world motion capture data [34], while the locations and rotations of object poses are randomly adjusted to simulate realistic motion [13]. Camera motion is controlled along predefined trajectories extracted from [21]. By collecting and animating these assets (characters, objects, and cameras), we generate 3D virtual scenes that closely mimic real-world dynamics. These scenes are then rendered with the Blender engine, producing image sequences $\{I_{1},I_{2},\cdots,I_{N}\}$, in which characters, objects, and the camera all move simultaneously. To maximize the diversity of the generated images, we randomly introduce HDR environments with various materials, textures, and lighting conditions, similar to those in [48].
@@ 3.2 Diffusion Model for Image Alignment
Dynamics-aware Mask Producing: Notably, our approach explicitly leverages the motion foreground mask to guide the network $\theta$ in learning denoising process. This guidance makes the network focus on the moving foreground objects, enabling the better reconstruction of aligned images with enhanced local details. The model relies on our DMP to generate motion foreground masks during training. Specifically, we first compute the correlation $\mathcal{C}(V_{1},V_{2})$ between $V_{1}$ and $V_{2}$ to extract sufficient motion information, then input $\mathcal{C}(V_{1},V_{2})$ into the mask predictor $\mathcal{M}$ to derive the predicted mask $\tilde{M}_{pred}$, as shown in Eq. 6:
$\mathcal{C}(V_{1},V_{2})$ $=\sum_{o\in\Theta}\left\langle V_{1}(u),V_{2}(u+o)\right\rangle,$ (6)
$\tilde{M}_{pred}$ $=\mathcal{M}(\mathcal{C}(V_{1},V_{2})),$
where $u$ represents the spatial coordinates of $V_{1}$, $\Theta$ denotes the neighborhood of coordinate $u$, and $o$ is the coordinate offset. Importantly, when using the foreground mask to guide the denoising process, we also aim to ensure that network notices the transition regions between foreground and background. To achieve smooth transitions at the edges of foreground objects, we apply the dilation operation $d_{r}$ to expand the produced mask $r$ times.
$V_{M}=V_{2}\times d_{r}(\tilde{M}_{pred})+V_{1}\times\{1-d_{r}(\tilde{M}_{pred})\},$ (7)
as shown in Eq. 7, the expanded foreground mask $d_{r}(\tilde{M}_{pred})$ is used to fuse the background of $V_{1}$ with the foreground of $V_{2}$, creating the mixed latent representation $V_{M}$, which serves as a condition to guide the diffusion network in generating the aligned image output.
Conditional Learning: As shown in Fig. 3, we concatenate $V_{1}$, $V_{2}$ and $V_{M}$ along the channel dimension to form the condition $x_{cond}$, then randomly select $x_{t}$ from the noise representations $\{x_{1},x_{2},\cdots,x_{T}\}$, allowing the diffusion model $\theta$ to learn the denoising process. This enables $\theta$ to handle any $x_{t}$ at an arbitrary timestep $t$, producing a denoised output $\hat{x}_{0}$. Similar to approaches in [45, 18, 28], we use the “predict $x_{0}$” denoising setting in our model. The condition $x_{cond}$ is concatenated with $x_{t}$ and the encoded timestep $t$ through positional embedding (PE) [47], and the result input is sent to the diffusion model to obtain the denoised result $\tilde{x}_{0}$:
@@
Table 1: Comparison of network inputs among DPFlow [36], COGS [17], GenWarp [40], and our DMAligner. Other optical flow networks (PWCNet [43], RAFT [44], FlowFormer++ [41], FlowDiffuser [28]) have the same input structure as DPFlow.
Method Main Input Structure
@@
Table 3: Quantitative comparison on Sintel and DAVIS dataset. Due to the lack of ground truth for precise evaluation, we compute LPIPS [64] and DreamSim [12] between $I_{pred}$ and $I_{1}$. Lower scores indicate better performance.
@@
where $\gamma$ is a variable weighting coefficient, and $d_{r}$ represents the dilation operation applied to the mask $r$ times. We also supervise the mask $\tilde{M}_{pred}$ using the cross-entropy loss, called as Mask Loss:
@@
In our experiments, the parameters $r$ (in $d_{r}$), $\lambda_{1}$, and $\lambda_{2}$ are set to 2, 2, and 0.1, respectively.
@@
After training, the diffusion model $\theta$ performs inference by denoising a standard Gaussian noise $x_{T}\in\mathcal{N}(0,1)$, conditioned on $V_{1}$, $V_{2}$ and $V_{M}$. This process reconstructs a reliable and accurate alignment result through recurrent denoising. Specifically, the model $\theta$ takes $x_{T}$, the condition, and the timestep $T$ as inputs according to Eq. 8, producing an initial coarse denoised output $\hat{x}_{0}^{(T)}$. Then, following the non-Markovian forward process [42] parametrized by $\sigma$ as shown in Eq. 13, it generates $x_{T-1}$ at timestep T1T-1. This process is repeated recurrently, producing a sequence of progressively refined denoised results $\hat{x}_{0}^{(T)},\hat{x}_{0}^{(T-1)},\cdots,\hat{x}_{0}^{(1)}$. Finally, $\hat{x}_{0}^{(1)}$ is passed through the decoder $\mathcal{D}$ to output the aligned image $I_{pred}$. The non-Markovian forward process is defined as:
@@ 4.1 Input Structure
As shown in Tab. 1, the input requirements for DPFlow [36], COGS [17], GenWarp [40], and our DMAligner are explained. Both DPFlow and DMAligner only require two consecutive image frames $I_{1}$ and $I_{2}$ for optical flow estimation and view synthesis, respectively. On the other hand, COGS relies on Gaussian splatting and requires not only the two input images $I_{1}$ and $I_{2}$, but also foreground masks ($M_{1},M_{2}$) and depth maps ($D_{1},D_{2}$), with the masks predicted by FC-CLIP [60] and depth estimated by Marigold [19]. These additional inputs are crucial for its 3D scene reconstruction. Although GenWarp is described as using only a single image $I_{2}$ as input, it actually relies on both $I_{1}$ and $I_{2}$ to obtain the relative camera motion $P_{I_{2}\to I_{1}}$, which can be derived from the optical flow $F_{I_{2}\to I_{1}}$ estimated by DPFlow, filtering out local motion to capture the overall camera movement [31, 32, 24]. Besides, it also needs camera intrinsics ($K$), depth ($D_{2}$) from Marigold. Therefore, GenWarp also relies on information from both $I_{1}$ and $I_{2}$. Unlike GenWarp and COGS, our DMAligner does not require additional geometric information such as camera intrinsics, depth maps, or pose transformations, making it a simpler and more direct approach for alignment.
@@ 4.3 Comparison with Existing Methods
Tab. 2 compares our DMAligner with some flow-warping methods and view-synthesis methods. In our DSIA dataset, all flow-warping methods are trained using rendered ground truth optical flow. We then align $I_{2}$ to $I_{gt}$ using back warping and compute PSNR and SSIM between $I_{pred}$ and $I_{gt}$. To address occlusion and ghosting artifacts from warping, these regions (determined by forward-backward consistency check [54]) are excluded from the metrics calculation, as marked by . Fig. 4 shows qualitative visual results on the DSIA test set, our DMAligner most closely matches the ground truth.
@@
We conduct qualitative comparisons on and the synthetic Sintel [3] dataset and the realistic DAVIS [4] dataset to evaluate the cross-domain performance of DMAligner and DSIA dataset. All competing methods and our DMAligner are trained on DSIA, and alignment results are directly tested on the Sintel Clean, Sintel Final training sets, and the DAVIS dataset. As there is no ground truth for precise evaluation, we calculate the feature distances between $I_{pred}$ and $I_{1}$ using LPIPS [64] and DreamSim [12]. Tab. 3 presents the quantitative results on Sintel and DAVIS, where DMAligner achieves the best average LPIPS and DreamSim scores. Fig. 6 shows examples from the DAVIS dataset, and Fig. 5 displays results from the Sintel test set. These results demonstrate that DMAligner delivers exceptional image alignment performance in both synthetic and real-world scenarios.
@@
In HDR imaging [53], LDR frames with alternating exposures (e.g., $L_{t-1}$$L_{t}$$L_{t+1}$) are aligned using flow warping before fusion. We replace this step with DMAligner, which aligns foreground and background more robustly, generating warped frames ($L_{t+1\to t}$, $L_{t-1\to t}$). Importantly, the fusion network and overall pipeline remain unchanged—only the aligned inputs differ. In Fig. 7, our method produces fewer artifacts, demonstrating strong generalization with existing HDR fusion network, which is expected to improve other downstream tasks as well.
Figure 8: Ablation experiments of DMP. The results of DMAligner without or with DMP are given, as well as the mask prediction $\hat{M}_{pred}$ and mask labels $M_{gt}$.
4.4 Ablation Study
In Tab. 4, we explore the impact of predict methods and DMP. Comparing (a)&(b) and (c)&(d)&(e), “predict $x_{0}$” is more effective for image alignment, as it directly targets image reconstruction. comparing (b)&(e), DMP resultes in more accurate alignment, as it allows the model to better focus on relevant areas. Fig. 8 also demonstrates that DMP enables the network to focus on the local dynamic foreground areas, improving the foreground details of the generated aligned images and achieving superior alignment.
Table 4: Ablation studies of predict methods and our DMP.
@@ 4.4 Ablation Study
Methods PSNR SSIM

### v2

@@ 1 Introduction
The alignment of consecutive frames is a critical task in computer vision, supporting a wide range of applications such as High-Dynamic-Range (HDR) imaging [49], image deblurring [52], burst image restoration [8], under-display camera imaging [7], video super-resolution [3], video denoising [55] and video stabilization [22]. Precise alignment ensures the accurate merging of information from consecutive frames, leading to enhanced image quality and better performance in subsequent tasks [35, 46]. Despite its significance, the task of image alignment remains challenging due to various factors, including scene dynamics, large camera motion, and varying lighting conditions.
@@
Figure 1: (a) Conventional image alignment based on optical flow and image warping, resulting in ghosting artifacts and occlusion. (b) Our DMAligner directly generates the complete alignment image via diffusion-based view synthesis.
@@
Recent advancements in deep learning have led to substantial progress in neural network-based alignment methods. Generally, image alignment can be divided into three levels: global-level, mid-level, and pixel-level alignment. Global-level alignment typically involves estimating a homography matrix, which can be accomplished using unsupervised deep learning frameworks [21]. This technique effectively aligns images under conditions of planar scenes or when the camera undergoes basic rotational or translational motion. Moreover, through global-to-local homography flow refinement, it can be naturally extended to local mesh-grid homography estimation [19], allowing it to overcome the limitations of a single homography [21]. However, global- and mid-level methods are restricted in their ability to address more complex scenarios, like parallax, non-rigid deformations, or significant foreground motion, since they assume single or multiple global transformations across the entire image.
@@
Another prominent technique involves the pixel-level approaches with optical flow and image warping [1]. Despite progress in image alignment, current optical flow-based methods face challenges that limit their effectiveness in practical applications. A major issue is the sensitivity of optical flow methods to brightness changes. Variations in lighting conditions, such as shadows, reflections, or changes in ambient light, can significantly distort the brightness constancy assumption, leading to incorrect motion estimation. Besides, the warping process inherent in optical flow-based alignment can introduce various artifacts. Inaccuracies in the motion field can lead to distortions such as stretching, compressing, or tearing of image regions. Additionally, the many-to-one mapping inherent in image warping on occlusion regions inevitably results in ghosting artifacts, as depicted in Fig. 1 (a), even with precise optical flow estimation. Such artifacts not only degrade the visual quality of the aligned frames but also affect the accuracy of subsequent processing tasks.
@@
In this paper, we introduce DMAligner, a diffusion-based framework for image alignment through alignment-oriented view synthesis. As depicted in Fig. 2, the desired image alignment result represents a scenario where the camera pose stays fixed at $P_{1}$ while time advances from $t_{1}$ to $t_{2}$. DMAligner is designed to handle the challenges in image alignment from a new perspective with a generation-based solution (see Fig. 1 (b)), exhibiting strong capabilities and avoiding issues associated with homography and flow-based image warping. Specifically, we first create Dynamic Scene Image Alignment (DSIA) dataset, designed to simulate typical image alignment challenges including camera shifts, moving objects, and fluctuating illumination. Moreover, we propose the Dynamics-aware Diffusion Training approach on the Latent Diffusion Model (LDM) [40] for learning conditional image generation. The key idea is to enhance the network’s capacity to recognize dynamic information during diffusion model training. This is realized through the Dynamics-aware Mask Producing (DMP) module, which effectively discriminates dynamic foreground regions from static backgrounds within the hidden features across frames. Additionally, we validate the effectiveness of DMAligner through extensive testing on synthetic datasets, including the proposed DSIA and Sintel [2], as well as on the widely-used real-world dataset DAVIS [36]. Both the quantitative and qualitative results demonstrate significant enhancements over contemporary methods. The contributions are summarized as follows:
@@ 2 Related Work
Traditional approaches primarily rely on feature-based methods, where keypoints are detected and matched across images, followed by geometric transformation estimation to align the images [18]. In the deep learning era, deep convolutional neural networks (DCNNs) have revolutionized image alignment in feature space [25]. Furthermore, the Transformer has also been utilized for feature-domain alignment, extending their application beyond standard RGB images [4] to directly process RAW image data [39, 37]. In image space, Global- and middle-level alignments typically involve estimating a homography matrix [21] and local mesh-grid homography [19], respectively. But they cannot handle complex scenarios, involving parallax, non-rigid deformations, or significant foreground motion. A notable advancement in image alignment is the adoption of optical flow-based image warping techniques. Optical flow calculates pixel-wise motion between consecutive images, enabling the warping of one image to align with another. With the advances of recent optical flow approaches [28, 6, 32], this method has found widespread use across a variety of video-based applications [26, 56]. Despite their success, warping in occlusion regions inevitably leads to ghosting artifacts, even with accurate flow, as shown in Fig. 1. In contrast, our DMAligner addresses the challenges in image alignment from a new perspective, performing alignment-oriented view synthesis using a diffusion-based framework.
@@
Diffusion models [11, 43, 40] own prominent generative performance for synthesis, and various approaches have been developed to improve the performance or efficiency. Different applications have shown good performances when empowered with diffusion models, including but not limited to, optical flow estimation [27], homography estimation [16], shadow removal [29], image rectangling [60], rolling shutter removal [51], HDR image reconstruction [24], low-light image enhancement [12, 13], raw image processing [23, 38], camouflaged object detection [58] and robot manipulation [5, 57, 47, 48]. In this work, we propose an end-to-end diffusion-based framework that specifically handles cross-frame dynamic information for image alignment.
@@ 3.1 DSIA Dataset Generation
Specifically, inspired by PointOdyssey [59] and Kubric [10], we assemble 25 characters and 100 objects with diverse shapes and materials from several open-source sources, including BlenderKit, Mixamo, GSO, and PartNet. These characters are animated using real-world motion capture data, while the locations and rotations of object poses are randomly adjusted to simulate real-world motion. Camera motion is controlled along predefined trajectories extracted from [17]. By collecting and animating these assets (characters, objects, and cameras), we generate 3D virtual scenes that closely mimic real-world dynamics. These scenes are then rendered with the Blender engine, producing image sequences $\{I_{1},I_{2},\cdots,I_{N}\}$, in which characters, objects, and the camera all move simultaneously. To maximize the diversity of the generated images, we randomly introduce HDR environments with various materials, textures, and lighting conditions.
@@ 3.2 Diffusion Model for Image Alignment
Dynamics-aware Mask Producing: To explicitly guide the denoising network to focus on moving foreground objects and enhance local detail reconstruction, we utilize a motion mask. First, we extract motion information by computing a normalized dot-product cost volume $\mathcal{C}$ between latents $V_{1}$ and $V_{2}$. For a spatial coordinate $\mathbf{u}$ and a displacement $\mathbf{d}$, the correlation is defined as:
$\mathcal{C}\{\mathbf{u},\mathbf{d}\}=\frac{1}{N}\sum_{k=1}^{N}V_{1}^{(k)}(\mathbf{u})\cdot V_{2}^{(k)}(\mathbf{u}+\mathbf{d}),$ (6)
where $N$ is the channel dimension. Assuming the input images $I_{1}$ and $I_{2}$ have a spatial resolution of $\mathbb{R}^{3\times H\times W}$ and the encoder $\mathcal{E}$ has a downsampling factor of $s$, the resulting latents $V_{1}$ and $V_{2}$ both belong to $\mathbb{R}^{N\times\frac{H}{s}\times\frac{W}{s}}$. Consequently, the correlation volume $\mathcal{C}\in\mathbb{R}^{D^{2}\times\frac{H}{s}\times\frac{W}{s}}$ (where $D$ represents the maximum displacement) is fed into a mask predictor $\mathcal{M}$ to generate the motion mask $\tilde{M}_{pred}=\mathcal{M}(\mathcal{C})$. To ensure smooth visual transitions at object boundaries, we expand this mask using a dilation operation $d_{r}$ (which is fully differentiable and implemented via the PyTorch “torch.nn.functional.max_pool2d” operation with a stride of 1). The dilated mask is subsequently used to fuse the background of $V_{1}$ with the foreground of $V_{2}$:
$V_{M}=V_{2}\odot d_{r}(\tilde{M}_{pred})+V_{1}\odot\{1-d_{r}(\tilde{M}_{pred})\}.$ (7)
This mixed latent $V_{M}$ serves as a critical conditioning signal to guide the diffusion network in generating the final aligned output.
Conditional Learning: As shown in Fig. 3, we concatenate $V_{1}$, $V_{2}$ and $V_{M}$ along the channel dimension to form the condition $x_{cond}$, then randomly select $x_{t}$ from the noise representations $\{x_{1},x_{2},\cdots,x_{T}\}$, allowing the diffusion model $\theta$ to learn the denoising process. This enables $\theta$ to handle any $x_{t}$ at an arbitrary timestep $t$, producing a denoised output $\hat{x}_{0}$. Similar to approaches in [27], we use the “predict $x_{0}$” denoising setting in our model. The condition $x_{cond}$ is concatenated with $x_{t}$ and the encoded timestep $t$ through positional embedding (PE), and the resulting tensor is sent to the diffusion model to obtain the denoised result $\tilde{x}_{0}$:
@@
Table 1: Comparison of network inputs among DPFlow [34], COGS [14], AccidentalGS [33], GenWarp [41] and our DMAligner. Other optical flow networks (PWCNet [44], RAFT [45], FlowFormer++ [42], FlowDiffuser [27]) have the same input structure as DPFlow.
Method Input Structure
@@
Table 3: Quantitative comparison on Sintel and DAVIS dataset. Due to the lack of ground truth for precise evaluation, we compute LPIPS and DreamSim [9] between $I_{pred}$ and $I_{1}$. Lower scores indicate better performance.
@@
where $\gamma$ is a variable weighting coefficient, and $d_{r}$ represents the dilation operation applied to the mask $r$ times. We also supervise the mask $\tilde{M}_{pred}$ using the cross-entropy loss, referred to as the Mask Loss:
@@
In our experiments, the downsampling factor of $s$, the parameters $r$ (in $d_{r}$), $\lambda_{1}$, and $\lambda_{2}$ are set to 4, 2, 2, and 0.1, respectively.
@@
After training, the diffusion model $\theta$ performs inference by denoising a standard Gaussian noise $x_{T}\in\mathcal{N}(0,1)$, conditioned on $V_{1}$, $V_{2}$ and $V_{M}$. This process reconstructs a reliable and accurate alignment result through recurrent denoising. Specifically, the model $\theta$ takes $x_{T}$, the condition, and the timestep $T$ as inputs according to Eq. (8), producing an initial coarse denoised output $\hat{x}_{0}^{(T)}$. Then, following the non-Markovian forward process [43] parametrized by $\sigma$ as shown in Eq. (13, it generates $x_{T-1}$ at timestep T1T-1. This process is repeated recurrently, producing a sequence of progressively refined denoised results $\hat{x}_{0}^{(T)},\hat{x}_{0}^{(T-1)},\cdots,\hat{x}_{0}^{(1)}$. Finally, $\hat{x}_{0}^{(1)}$ is passed through the decoder $\mathcal{D}$ to output the aligned image $I_{pred}$. The non-Markovian forward process is defined as:
@@ 4.1 Input Structure
As detailed in Tab. 1, our DMAligner and DPFlow [34] are highly efficient, requiring only two consecutive frames ($I_{1},I_{2}$) for alignment-oriented view synthesis and optical flow estimation, respectively. In contrast, 3DGS-based methods heavily depend on auxiliary priors for scene reconstruction. Beyond two input images ($I_{1}$ and $I_{2}$), COGS [14] requires depth maps ($D_{1},D_{2}$) estimated by Marigold [15] and foreground masks ($M_{1},M_{2}$) predicted by FC-CLIP [54]. AccidentalGS [33] similarly relies on two input images ($I_{1}$ and $I_{2}$), depth priors ($D_{1},D_{2}$) and camera intrinsics ($K$). Furthermore, while GenWarp [41] claims single-image input ($I_{2}$), it implicitly requires both $I_{1}$ and $I_{2}$ to estimate relative camera motion $P_{I_{2}\to I_{1}}$ via optical flow [30, 31, 20], alongside camera intrinsics ($K$) and depth ($D_{2}$) from Marigold [15]. Ultimately, unlike these geometry-dependent approaches, our DMAligner achieves accurate alignment without requiring depth, masks, intrinsics, or pose transformations, offering a much simpler and more direct framework.
@@ 4.3 Comparison with Existing Methods
Tab. 2 compares our DMAligner with some flow-warping methods and view-synthesis methods. In our DSIA dataset, all flow-warping methods are trained using rendered ground truth optical flow. We then align $I_{2}$ to $I_{gt}$ using back warping and compute PSNR and SSIM between $I_{pred}$ and $I_{gt}$. To address occlusion and ghosting artifacts from warping, these regions (determined by forward-backward consistency check [50]) are excluded from the metrics calculation, as marked by . Fig. 4 shows qualitative results on the DSIA test set, where our DMAligner most closely matches the ground truth.
@@
We conduct qualitative comparisons on the synthetic Sintel [2] dataset and the real-world DAVIS [36] dataset to evaluate the cross-domain performance of DMAligner and DSIA dataset. All competing methods and our DMAligner are trained on DSIA, and alignment results are directly tested on the Sintel Clean, Sintel Final training sets, and the DAVIS dataset. As there is no ground truth for precise evaluation, we calculate the feature distances between $I_{pred}$ and $I_{1}$ using LPIPS and DreamSim [9]. Tab. 3 presents the quantitative results on Sintel and DAVIS, where DMAligner achieves the best average LPIPS and DreamSim scores. Fig. 6 shows examples from the DAVIS, and Fig. 5 displays results from the Sintel. These results demonstrate that DMAligner delivers exceptional image alignment performance in both synthetic and real-world scenarios.
@@
In HDR imaging [49], LDR frames with alternating exposures (e.g., $L_{t-1}$$L_{t}$$L_{t+1}$) are aligned using flow warping before fusion. We replace this step with DMAligner, which aligns foreground and background more robustly, generating warped frames ($L_{t+1\to t}$, $L_{t-1\to t}$). Importantly, the fusion network and overall pipeline remain unchanged—only the aligned inputs differ. In Fig. 7, our method produces fewer artifacts, demonstrating strong generalization with existing HDR fusion networks.
Table 4: Ablation studies of predict targets and our DMP.
@@
Targets PSNR SSIM
@@
Ablation Study. Tab. 4 ablates prediction targets and DMP. Experiments (a)-(e) show that “predict $x_{0}$” achieves better alignment by directly targeting image reconstruction. Besides, (b) vs. (e) demonstrates that DMP enhances alignment by focusing the model on relevant dynamic regions.
Robustness Analysis. Stress tests in Fig. 8 reveal that DMAligner maintains a slower degradation rate and outperforms others even under large motion magnitude gaps.